How do I calculate the Ksp, Delta G, and Delta S from my data table? The experiment aims to find the solubility curve of potassium nitrate dissolving in water. For example, when 2 grams of KNO3 dissolve in 5 ml of H20, crystals reappeared at 25° celsius.


1
Expert's answer
2022-05-04T03:28:02-0400

KNO3 (s) ⇌ K+ (aq) + NO3- (aq) 


 Calculate the concentration of KNO3 

2.0 g × (1 𝑚𝑜𝑙 / 101.1 𝑔) × (1/1 𝐿) = 0.02 mol/L 


Calculate the Ksp 

Ksp = [K+][NO3-

[K+] - molar concentration of K+ (mol/L), [NO3-] - molar concentration of NO3- (mol/L) 

Ksp = 0.02*0.02 = 0.0004


Determine the ΔG based on the Ksp 

ΔG = -RTlnKsp    

 R = 8.314 J/L*mol, T - temperature in Kelvin 

ΔG = -(8.314 J/K*mol)(25°C + 273 K)(ln0.0004) = 19374.61 J/mol 


ΔG = ΔH – TΔS 

The ΔH and ΔS based on data at different temperatures

a plot of lnKsp versus 1/T yields a straight line with a slope of –ΔH/R and a y-intercept of ΔS/R. So the ΔH and ΔS can be calculated based on the linear equation obtained from the linear fitting of the graph.